Circular No. 7517

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2000ej IN IC 1371 M. Schwartz, Cottage Grove, OR; and W. D. Li and A. V. Filippenko,

University of California at Berkeley, report the discovery of an apparent supernova (mag ∼ 18) on unfiltered images taken with the 0.5-m Tenagra III telescope on Nov. 2.2 and 6.2 UT in the course of the LOTOSS (cf. IAUC 7514). The new object is located at α = 21^{h}20^{m}15^{s}.66, δ = −4^{o}52′40′′.5 (equinox 2000.0), which is 5′′.8 south of the nucleus of IC 1371. Photometry of the supernova is difficult because of its proximity to the bright nucleus of the host galaxy. An unfiltered image taken with the 0.8-m Katzman Automatic Imaging Telescope on Sept. 29.2 showed nothing at the position of SN 2000ej (limiting mag ∼ 19.0).

SUPERNOVA 2000dw IN UGC 11955 Filippenko and Li also report that inspection of noisy CCD spectra

(range 330–1000 nm) of SN 2000dw (cf. IAUC 7511), obtained on Nov. 1 UT with the Shane 3-m reflector at Lick Observatory, reveals that the object is a normal type-II supernova. Hα has a classic P-Cyg profile.

SUNGRAZING COMETS Further to IAUC 7514 and 7485, D. Hammer reports his measurements

for eight additional Kreutz sungrazing comets found on SOHO website images. Comet C/1997 T7 was found by X. Leprette; the other seven objects were found by M. Oates. All eight comets were seen in the C2 coronagraph, but only only C/1999 V2 was also visible with the C3 instrument. The reduced observations and orbital elements by B. G. Marsden are given on the MPECs cited below.

Comet  1997 UT  α_{2000}  δ_{2000}  MPEC
C/1997 T6 Oct. 1.994  12^{h}23.^{m}2  − 3^{o}58′  2000-V14
C/1997 T7  8.335  12 46.5  − 6 30  2000-V22
Comet  1999 UT  α_{2000}  δ_{2000}  MPEC
C/1999 U7 Oct. 27.129  13^{h}57.^{m}0  −13^{o}32′  2000-V14
C/1999 U8  29.326  14 05.9  −14 21  2000-V14
C/1999 U9  30.271  14 10.1  −14 44  2000-V14
C/1999 V2 Nov. 7.721  14 37.8  −18 29  2000-V22
C/1999 V3  14.746  15 15.3  −19 44  2000-V22
C/1999 V4  15.538  15 18.6  −19 54  2000-V22
2000 November 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green